Non-exact equation and Integrating Factor


Given the equation below, solve for the general solutions.


w = dependent variable, v = independent variable


1. w(2v - w + 1) dv + v (3v - 4w + 3) dw = 0


1
Expert's answer
2021-06-23T17:23:47-0400
"P(v, w)=2vw-w^2+w"

"P_w=2v-2w+1"


"Q(v,w)=3v^2-4vw+3v"

"Q_v=6v-4w+3"

"P_w-Q_v=2v-2w+1-6v+4w-3"

"=-4v+2w-2"

"\dfrac{P_w-Q_v}{P}=\dfrac{-4v+2w-2}{2vw-w^2+w}=-\dfrac{2}{w}=-\psi(w)"


"\mu(w)=e^{\int{2 \over w}dw}=w^2"


"(2vw^3-w^4+w^3)dv+(3v^2w^2-4vw^3+3vw^2)dw=0"

"M_w=6vw^2-4w^3+3w^2"


"N_v=6vw^2-4w^3+3w^2"

"M_w=6vw^2-4w^3+3w^2=N_v"

Then we write the system of two differential equations that define the function "u(v,w):"


"u_v=M(v,w)"

"u_w=N(v,w)"


"u(v,w)=\int(2vw^3-w^4+w^3)dv+\varphi(w)"

"=v^2w^3-vw^4+vw^3+\varphi(w)"

"u_w=N(v,w)"

"u_w=3v^2w^2-4vw^3+3vw^2+\varphi'(w)"

"=3v^2w^2-4vw^3+3vw^2"

Then

"\varphi'(w)=0"

"\varphi(w)=C_1"

The general solution of the differential equation


"w(2v-w+1)dv+v(3v-4w+3)dw=0"

is given by


"v^2w^3-vw^4+vw^3=C"